# Lasso Multinomial Performance Indicators for in-play Basketball Data

### v1

@@ Abstract
A typical approach to quantify the contribution of each player in basketball uses the plus/minus approach. Such plus/minus ratings are estimated using simple regression models and their regularised variants with response variable either the points scored or the point differences. To capture more precisely the effect of each player and the combined effects of specific lineups, more detailed possession-based play-by-play data are needed. This is the direction we take in this article, in which we investigate the performance of regularized adjusted plus/minus (RAPM) indicators estimated by different regularized models having as a response the number of points scored in each possession. Therefore, we use possession play-by-play data from all NBA games for the season 2021-22 (322,852 possessions). We initially present simple regression model-based indices starting from the implementation of ridge regression which is the standard technique in the relevant literature. We proceed with the lasso approach which has specific advantages and better performance than ridge regression when compared with selected objective validation criteria. Then, we implement regularized binary and multinomial logistic regression models to obtain more accurate performance indicators since the response is a discrete variable taking values mainly from zero to three. Our final proposal is an improved RAPM measure which is based on the expected points of a multinomial logistic regression model where each player’s contribution is weighted by his participation in the team’s possessions. The proposed indicator, called weighted expected points (wEPTS), outperforms all other RAPM measures we investigate in this study.
Keywords: plus/minus, RAPM, regularization, logistic regression, expected points (EPTS)
@@ 1 Introduction
Managing the workload of players is a complex process that challenges basketball coaches and analysts worldwide. This is one of the factors that account for the popularity of players’ evaluation criteria. Particularly in team sports, one of the most relevant and valuable metrics is the impact a player has on the outcome of his team. As the renowned Dean Oliver stated “Teamwork is the element of basketball most difficult to capture in any quantitative sense” [7].
A common way to assess the performance of teams and players is to use plus/minus indices. WINVAL is a software system that implements the adjusted plus/minus technique for basketball, which was created by Wayne Winston and Jeff Sagarin in 2002 [11]. It calculates the average point difference for each NBA player’s team per possession when the player is on the court. This number indicates how well the player’s team performed when he was playing. However, this technique has a major flaw: it favors players who often play with good teammates or against weak opponents. These players tend to appear more effective than they actually are.
Using only the information on the players’ teammates and opponents on the court, [9] and, later, [5, 6] suggested a ridge regression approach to measure the players’ contribution. The index was named Regularized Adjusted Plus/Minus (RAPM). The model has become popular and it was widely used in players’ performance analysis because of its benefits: i) it produces low-variance estimates; ii) it provides an evaluation metric for all players; iii) it provides more stable evaluation metrics for players that frequently play together because it handles multicollinearity in the fitted model; and iv) it allows easy interpretation of the RAPM evaluation metrics via the underlying fitted linear regression.
@@
Our initial intuition that the multinomial model is more suitable for modelling the points per possession and for producing player evaluation metrics is confirmed since the obtained wEPTS evaluation index is found to be superior according to selected external validation criteria. Moreover, this metric seems to partially solve the problem related to the overestimation of the performance of low-time players. Another characteristic of the proposed regularized multinomial approach is that the overall EPTS (and wEPTS) index can be decomposed in three different RAPM measures, one for each possible scoring outcome (one, two and 3+). Therefore, the method can spot the best-performed players according to each scoring outcome. For example, some players might be more influential by participating in two-pointer offences and other players might participate more when their team scores a three-pointer. Our method will identify both of these types of players and account for their contribution. This is a unique characteristic of our approach that no other previously proposed method has.
@@
The structure of the paper is as follows. Section 2 describes the data and the methodology used in this study. Section 3 presents the results of our analysis regarding RAPM and EPTS ratings and, in Section 4, we introduce the finally proposed weighted expected points (wEPTS) rating. Section 5 concludes with some interesting findings and discussions.
@@ 2.1 Initial Raw Dataset
In this study, play-by-play possession NBA data were used from the 2021-2022 season. Data were collected from NBA logs. The data comprised of possessions as the study unit of analysis (i.e. observation) and the related information (features) of each possession, such as points scored and playing lineups of the two opponent teams. The original (unprocessed) dataset contained 322,852 observations (possessions) and 13 variables. A description of the available variables follows:
@@
    O1–O5 and D1–D5: String variables that refer to the players of the team in possession (i.e. in offence) and the opposing team (i.e. in defence), respectively, in each possession. Each string/character comprises by the team’s (abreviated) name & player’s name, e.g., LAC22 Ivica-Zubac.
@@ 2.2 Final Dataset
The original raw dataset was approrpiately transformed in order to be used in the subsequent analysis. Specifically, the players’-related variables (O1–O5 and D1–D5) were transformed into a series of dichotomous variables (one offensive and one defensive for each player who participated in at least one game of the NBA league). These dichotomous variables denote the presence or absence of each player in the field, in each possession, with the team in possession or the team in defence. For this dataset, we have considered a total of 717 players. Thus, the final data matrix is an $n\times m$ sparse matrix with $n=322,852$ rows (possessions) and $m=1,437$ columns with the features used in the analysis. The total set of features consists of $717\times 2$ binary variables used to evaluate the offensive and defensive contribution of each player (i.e. each player is split into their offensive and defensive coefficients) plus three extra columns from the original raw dataset: home_off (binary), season_type (factor with 2 levels), and pts (points per possession).
@@ 2.3 Models and Analysis
In this section, we start from the standard approach in the Basketball bibliography for obtaining RAPM measures which is based on the normal regression model using ridge regression. The main drawback of such a model is that the normal distribution attached to it is incompatible with the use of the number of points scored in each possession as the dependent efficiency variable. Therefore, we proceed with models that are more suitable for this kind of response. Initially, we use a binomial logistic regression model with the binary indicator of scoring or not scoring per possession as a dependent efficiency variable. Although this model is more appropriate mathematically than the normal regression model, there is a loss of information since all different types of scoring are merged into one unified category. This loss of information (i.e. the different types of scores) is quite important for describing the nature of Basketball. For this reason, we proceed by proposing a multinomial regression logistic regression model which takes into consideration the three different outcomes of a possession. In all the above models, additionally to the ridge approach, we have also implemented lasso which has the characteristic that places specific players in an “average” category. Overall, lasso seems to outperform the ridge approach in many different perspectives that will be later explained in Section 3.
@@ 2.3.1 Low-time players
Prior to proceeding with the implemented models, we need to introduce the concept of Low-time players (LTPs). In the following, we use this term for the players who participate for less than a threshold of minutes per game, usually due to their restricted role in the team. Their performance may affect the evaluation metrics that are widely used to assess the contribution of basketball players to their teams’ success. For instance, conventional evaluation metrics do not take into consideration the context and quality of their playing time. Moreover, such metrics heavily depend on the performance of the other players on the court, as well as the quality of the opposing team [^1] https://www.nbastuffer.com/analytics101/plus-minus/. On this perspective, [4] argues that the quality and the performance of the team that a low-time player participates greatly affects their plus/minus ratings. As a result, poor players on good teams are overrated, while good players on poor teams are underrated.
A common way to resolve the problem is to impose a threshold on the playing time and exclude the corresponding LTPs from the calculation of the performance evaluation metrics; see for example in [9] and [6].
Subsequently, in the following, we consider as LTPs all players with playing time less than 200 minutes for the whole of the season. This threshold identifies 232 (out of 717) as LTPs which is the $32\%$ of the players in our dataset.
@@ 2.3.2 Linear normal regression with shrinkage methods
for $i=1,\dots,n$; where $n$ the total number of possessions in our dataset, $K$ is the total number of players under consideration, $Pts_{i}$ is the number of points achieved in possession $i$, $\mu_{i}$ is the expected points for possession $i$ for this model, $X_{ik}^{o}$ and $X_{ik}^{d}$ are the indicators of player $k$ in possession $i$ for the offensive and the defensive team, $X_{ij}$ refers to the extra covariate $j$ for possession $i$, $p$ is the number of extra covariates used in the model formulation, $\beta^{o}_{k}$ and $\beta^{d}_{k}$ are the offensive and defensive, respectively, RAPM coefficients for player $k$, $b_{k}$ are the effects of any extra covariates and $b_{0}$ is the overall constant of the model (usually referring to a reference average player obtained by players with zero or very close to zero RAPM coefficients). Finally, $\sigma^{2}$ is the error (or unexplained) variance which specifies the accuracy of the fitted model. In many published articles [see for example in 5, 6], the number of points per possession is multiplied by a factor of 100.
@@
Hence interest lies in the estimation of $\beta_{k}^{o}$ and $\beta_{k}^{d}$ coefficients (i.e. on ORAPM and DRAPM, respectively). The constant term $b_{0}$ here is a nuisance parameter with no direct or meaningful interpretation under the current formulation. The RAPM coefficients $\beta^{o}_{k}$ and $\beta^{d}_{k}$ measure the offensive and defensive, respectively, contribution of player $k$ for $k=1,\ldots,K$ players. Pairwise comparison between two players (denoted by $k$ and $m$) can be achieved by considering the RAPM differences $\beta^{o}_{k}-\beta^{o}_{m}$, $\forall k,m=1,\ldots,K,k\neq m$. These differences represent the mean discrepancy in the expected points scored by each player’s team when player $k$ substitutes player $m$ in a specific possession, with all other players on the court staying the same. Similarly, the difference between two defensive RAPMs $\beta^{d}_{k}-\beta^{d}_{m}$ ($\forall k,m=1,\ldots,K,k\neq m$) expresses the average discrepancy in points of each player’s team conceded (by the opponent team) when players $m$ plays instead of player $k$ with all other players on the court staying the same. Hence, the larger the difference is, the higher the offensive and defensive contribution of $k$ player.
@@
The related literature mostly suggests ridge regression to derive RAPM ratings. Ridge regression shrinks coefficients towards zero but does not set them exactly equal to zero. This has the advantage that every player will get a non-zero evaluation coefficient for his offensive and defensive contributions; for more details about the ridge regression see [3] and in [9, 5, 4, 2] for the implementation in obtaining basketball RAPMs evaluation metrics.
In the implementation of ridge regression, for the estimation of RAPM coefficients, we have used 10-fold cross-validation in order to specify the tuning shrinkage parameter of ridge regression (often denoted by $\lambda$). In the ridge bibliography, one of the standard choices is to use $\lambda$ which minimizes the average mean root square error (RMSE) obtained from cross-validation. This value was found to be equal to $\lambda_{min}=241.07$ which is considerably lower than the standard value of 2000 reported in RAPM bibliography [see for example in 10]. The shrinkage percentage with the choice of $\lambda_{min}$ was extremely high ($>99\%$).
A clear issue that emerged in the RAMP-based player rankings, is that low-time players (LTPs), namely players with less than 200 minutes played in the entire season, appear at the top of the evaluation list with exceedingly high RAPM ratings. This is in line with the literature; refer to for instance in [7]. To address this issue, we opted to implement lasso instead of ridge and, in second step, to exclude LTPs from the analysis or to eliminate their RAPMs from the regression evaluation metric.
@@
Lasso has the extra property, in comparison to ridge, that shrinks smaller coefficients exactly to zero. This provides a more meaningful interpretation of the intercept coefficient of the implemented model, since now it represents the average contribution of a reference player in each possession. By reference player, we mean any player belonging in the group of players with zero coefficients. A total of 553 offensive and 574 defensive players (out of the 717) have been identified to belong in this reference group. Therefore 184 offensive and 143 defensive players were found with non-zero lasso RAPM ratings. To derive the lasso RAPM ratings we have used the shrinkage parameter value of $\lambda_{min}=0.27$ obtained by a 10-fold cross-validation.
@@
Figure 1 presents the comparison between the ridge and the lasso RAPMs. From this comparison, it is evident that the ridge coefficients differ considerably from the lasso ones. The main difference lies in the zero-coefficients of lasso. However, when zero coefficients are excluded, a strong linear relationship emerges between the lasso and ridge offensive RAPMs with Pearson’s linear correlation coefficient equal to $78.1\%$ ($59.6\%$ when including the zero values of lasso). The corresponding relationship between defensive RAPMs is lower with correlation equal to $61.8\%$ ($55.2\%$ when including the zero values of lasso).
Figure 1: Ridge vs. lasso regression RAPM ratings (considering all players)

### v2

@@ Abstract
A typical approach to quantify the contribution of each player in basketball uses the plus-minus method. The ratings obtained by such a method are estimated using simple regression models and their regularized variants, with response variable being either the points scored or the point differences. To capture more precisely the effect of each player, detailed possession-based play-by-play data may be used. This is the direction we take in this article, in which we investigate the performance of regularized adjusted plus-minus (RAPM) indicators estimated by different regularized models having as a response the number of points scored in each possession. Therefore, we use possession play-by-play data from all NBA games for the season 2021-22 (322,852 possessions). We initially present simple regression model-based indices starting from the implementation of ridge regression which is the standard technique in the relevant literature. We proceed with the lasso approach which has specific advantages and better performance than ridge regression when compared with selected objective validation criteria. Then, we implement regularized binary and multinomial logistic regression models to obtain more accurate performance indicators since the response is a discrete variable taking values mainly from zero to three. Our final proposal is an improved RAPM measure which is based on the expected points of a multinomial logistic regression model where each player’s contribution is weighted by his participation in the team’s possessions. The proposed indicator, called weighted expected points (wEPTS), outperforms all other RAPM measures we investigate in this study.
Keywords: plus-minus, RAPM, regularization, logistic regression, expected points (EPTS)
@@ 1 Introduction
In team sports, one of the most relevant and valuable metrics is the impact a player has on the outcome of his team. As the renowned Dean Oliver stated “Teamwork is the element of basketball most difficult to capture in any quantitative sense” [21]. Using only the information of the players’ teammates and opponents on the court, [23] introduced Adjusted Plus-Minus (APM) to measure the players’ contribution. Later, [16, 17] suggested a ridge regression approach to address multicollinearity issues in APM, leading to the development of Regularized Adjusted Plus-Minus (RAPM).
@@
Our initial intuition that the multinomial model is more suitable for modelling the points per possession and for producing player evaluation metrics is confirmed since the obtained wEPTS evaluation index is found to be superior according to selected external validation criteria. Moreover, this metric seems to partially solve the problem related to the overestimation of the performance of low-time players. Another characteristic of the proposed regularized multinomial approach is that the overall EPTS (and wEPTS) index can be decomposed in three different RAPM measures, one for each possible scoring outcome (one, two and 3+). Therefore, the method can spot the best-performed players according to each scoring outcome. For example, in the modern NBA the value of the three point shot cannot be overstated. However, there can be players that while they do not take a high volume of threes they contribute significantly to their teammates taking and making higher quality threes. For example, Giannis Antetokounmpo and LeBron James are not high volume three point shooters, but their playmaking abilities, in conjunction with their gravity allows them to generate high quality three point shots for their teammates. Our model will be able to identify these contributions from players and this is a unique characteristic of our approach that no other previously proposed method has.
@@
The structure of the paper is as follows. Section 2 presents the main background in sports and basketball analytics regarding plus-minus, Section 3 describes the data and the methodology used in this study. Section 4 presents the results of our analysis regarding RAPM and EPTS ratings and, in Section 5, we introduce the finally proposed weighted expected points (wEPTS) rating. Section 6 concludes with some interesting findings and discussions.
2 Background
Due to developments in technology and the growing availability of detailed game data, the application of advanced analytics in basketball has experienced significant growth in recent years. Initially, basketball analytics concentrated on fundamental statistical metrics, such as points per game and field goal percentage but, in recent years, more advanced and sophisticated methods have been developed. As [21], discusses in Basketball on Paper, understanding basketball from an analytical perspective involves considering not only individual metrics but also team dynamics and the interaction between offensive and defensive strategies.
In the context of player evaluation, the primary challenge lies in “isolating” the contribution of individual players from their teammates and opponents. This led to the development of various metrics designed to quantify a player’s impact on their team. Plus-minus ratings have been a key tool in evaluating player performance in team sports, especially in basketball. Traditional plus-minus assesses a team’s net performance, typically expressed as a point differential over 100 possessions, when a specific player is on the court. However, its simplicity can be misleading; players on strong teams who face weaker opponents may appear to have excessive ratings, while elite players on underperforming teams might receive an unexpectedly low rating that does not correspond to their quality of play. The introduction of Adjusted Plus-Minus (APM) by [23] is an important advance over traditional plus-minus by accounting for both teammate and opponent quality, reducing the biases of earlier models. While the introduction of APM was well received by both academics and professional basketball analysts for capturing players’ impact without needing to assign in-game actions or box-score statistics, it has certain limitations; (a) its instability, (b) the presence of extreme outliers and (c) collinearity of player indicators. Regarding instability, a player’s APM can vary greatly from season to season, making this metric unreliable for systematic evaluation. Concerning the presence of extreme outliers, this refers to certain players who may exhibit unexpectedly high or low APM scores. Additionally, we have the issue of low-minute players, who appear as outliers primarily due to two reasons: the lack of playing data and the quality of their data, as they usually play during time periods that are not critical for their team (garbage time). Finally, collinearity refers to the fact that specific players often play together in the same lineup, making difficult to estimate each player’s individual contribution.
Rosenbaum’s [2004] regression-based APMs have inspired many researchers in the field of sports analytics. Starting from basketball, Ilardi and Barzilai [16, 17] implemented the model by incorporating ridge regression, a regularization technique that resolves multicollinearity issues. Ridge regression is the main tool for the development of the Regularized Adjusted Plus-Minus (RAPM) model, as it provides stable estimates, especially for players with limited playing time. RAPM’s advantages include: i)low-variance estimates; ii)an evaluation metric for all players; iii)more stable evaluation metrics for players that frequently play together because it handles multicollinearity in the fitted model; and iv)easy interpretation of the RAPM evaluation metrics via the underlying fitted linear regression.
Another modification was proposed by [18], who developed two weighted least squares regression models to estimate a National Hockey League (NHL) player’s effect on his team’s success in scoring and preventing goals, independent of that player’s teammates and opponents. One year later, [19] additionally applied the ridge regularization technique in his implementation.
A notable advancement in the field was the development of WINVAL software system, introduced by [29]. Unlike traditional methods that focus on the average point difference per possession, WINVAL evaluates the impact a player has on probability of win through the score difference when he is on the court. However, the system still faces limitations, such as over-evaluating players who consistently play alongside strong teammates or face weaker opponents at meaningless points of the game (like garbage time).
One of the main characteristics of RAPM estimated via ridge regression is that the method assigns non-zero ratings to all players. In contrast, lasso regression (Least Absolute Shrinkage and Selection Operator), introduced by Tibshirani [28], treats RAPM ratings differently by shrinking the ratings of specific players to zero. In sports analytics, this regularization method is not so popular. However, [7], and, later, [8] presented Bayesian lasso in logistic regression to model the probability of scoring a goal by the home team in hockey by taking advantage of the variable selection implied by lasso.
Following the trend of model-based plus-minus ratings, weighted metrics are another alternative approach to measuring player performance. For instance, [20] noted that APM does not have the same impact in soccer as it does in basketball and hockey, since soccer is a low-scoring sport, with a low number of substitutions. Therefore, they developed the Augmented APM by combining FIFA ratings with APM, resulting in stronger predictive ability than standard APM or a simple regression model using only FIFA ratings. This metric, also, considers the problem of collinearity among players. The idea is to develop APM through a Bayesian framework and incorporate FIFA ratings into the prior distribution. [24] introduced the use of a different combination of features in the model used for the estimation of RAPMs for individual soccer player performance called the weighted plus-minus metric. This metric is also based on considering the opponent team’s strength (according to betting odds) and the goal’s importance (considering the impact of a goal on the game’s outcome). Finally, feature-enhanced ratings have also been popular in basketball in recent years. They are mainly based on box-score statistics and tracking data; see for example RAPTOR PIE, DARKO and LEBRON; see in NBAstuffer website for a detailed list including references (https://www.nbastuffer.com/analytics101/).
Bayesian models in basketball analytics have gained popularity the recent years. [4] introduced an adjusted plus-minus model within a Bayesian framework to estimate offensive and defensive ratings, assuming these ratings follow distinct Gaussian distributions. They incorporated player rating variations over time by using each season’s ratings on the prior season’s results, adjusted toward a prior mean. Later, [3] extended this idea by applying Bayesian linear regression with independent Laplacian priors for player and team components, suggesting the ratio of the posterior mean to the posterior standard deviation of the player’s partial effect to compare players. A recent noteworthy contribution comes from [9], who applied Bayesian approaches to improve RAPM by incorporating prior information about players, leading to more stable estimates in the presence of sparse or noisy data. Moreover, they have introduced the estimation of lineup effects as an additional factor (positive or negative) for each player, depending on the lineup in which they are included.
Regarding the football (soccer) players’ evaluation, a broad variety of research works exists. In the field of football player performance evaluation, the contribution of Professor Hvattum is notable. He initially published a plus-minus methodology in football [26] — adjusted plus-minus ratings obtained by a ridge multiple linear regression model — which has been extended by suggesting an additional component in each player rating corresponding to the league they participate [27]. He also created and maintains a YouTube channel called “Football Player Ratings” [^1] https://www.youtube.com/@footballplayerratings, where he presents and discusses statistical models for evaluating individual football players in a straightforward and accessible manner through clear, easy-to-understand videos.
In a later work, [14] proposed an alternative plus-minus rating model, and introduced individual offensive and defensive contribution football players. By considering segments of a football match with fixed lineup, ratings are determined by minimizing the squared difference between the actual goals scored and the expected goals based on players’ ratings, accounting for effects like home-team advantage, segment duration, red card impact, age effect, and league strength. Consequently, the numerical ratings represent each player’s contribution to goals scored per 90 minutes. He, also, applied ridge regularization technique to obtain more stable ratings [15], while he proposed weights for each segment difference, depending on the timepoint of the segment, the duration of the segment, and the game states.
Despite the clear diversity of plus-minus ratings in performance evaluation, there remains a gap regarding the lack of use of logistic models in the field — particularly within basketball analytics — presenting an intriguing area for further investigation. Consequently, based on the literature, the primary aim of this study is to develop multinomial logistic models for play-by-play data, inspired by the concept of Regularized Adjusted Plus-Minus (RAPM), by applying lasso as an alternative regularization technique and by considering a new weighted evaluation metric based on each player’s participation in the game.
@@ 3.1 Initial Raw Dataset
In this study, play-by-play possession NBA data were used from the 2021-2022 season. Data were collected from NBA logs. The unit of study (i.e., observation) is each possession, for which points scored and the lineups of the two opposing teams are collected as features. The original (unprocessed) dataset contained 322,852 observations (possessions) and 13 variables. A description of the available variables follows:
@@
    O1–O5 and D1–D5: String variables refer to the players (and their teams) participating in each possession. Each string consists of the team’s abbreviated name and the player’s name, e.g., LAC22 Ivica Zubac. Features O1–O5 refer to the players of the team in possession (i.e., on offense), while D1–D5 refer to the players of the opposing team (i.e., on defense).
@@ 3.2 Final Dataset
The original raw dataset was appropriately transformed in order to be used in the subsequent analysis. Specifically, the players’-related variables (O1–O5 and D1–D5) were transformed into a series of dichotomous variables (one offensive and one defensive for each player who participated in at least one game of the NBA league). These dichotomous variables denote the presence or absence of each player in the field, in each possession, with the team in possession or the team in defence. For this dataset, we have considered a total of 717 players. Thus, the final data matrix is an $n\times m$ sparse matrix with $n=322,852$ rows (possessions) and $m=1,437$ columns with the features used in the analysis. The total set of features consists of $717\times 2$ binary variables used to evaluate the offensive and defensive contribution of each player (i.e. each player is split into their offensive and defensive coefficients) plus three extra columns from the original raw dataset: home_off (binary), season_type (factor with 2 levels), and pts (points per possession).
@@ 3.3 Models and Analysis
In this section, we start from the standard approach in the basketball bibliography for obtaining RAPM measures which is based on the normal regression model using ridge regression. The main drawback of such a model is that the normal distribution attached to it is incompatible with the use of the number of points scored in each possession as the dependent efficiency variable. Therefore, we proceed with models that are more suitable for this kind of response. Initially, we use a binomial logistic regression model with the binary indicator of scoring or not scoring per possession as a dependent efficiency variable. Although this model is more appropriate mathematically than the normal regression model, there is a loss of information since all different types of scoring are merged into one unified category. This loss of information (i.e. the different types of scores) is quite important for describing the nature of basketball. For this reason, we proceed by proposing a multinomial regression logistic regression model which takes into consideration the three different outcomes of a possession.
In all the above models, additionally to the ridge approach, we have also implemented lasso which has the characteristic that places specific players in an “average” category. Overall, lasso seems to outperform the ridge approach in many different perspectives that will be later explained in Section 4.
Finally, considering our reviewers’ suggestions, we have also included in our comparisons a Poisson regression and an ordinal logistic model. The results of these models confirmed our initial intuition that such models are not appropriate for the specific case under study.
@@ 3.3.1 Low-time players
Prior to proceeding with the implemented models, we need to introduce the concept of Low-time players (LTPs). In the following, we use this term for the players who participate for less than a threshold of minutes per game, usually due to their restricted role in the team. Their performance may affect the evaluation metrics that are widely used to assess the contribution of basketball players to their teams’ success. For instance, conventional evaluation metrics do not take into consideration the context and quality of their playing time. Moreover, such metrics heavily depend on the performance of the other players on the court, as well as the quality of the opposing team [^2] https://www.nbastuffer.com/analytics101/plus-minus/. From this perspective, [13] argues that the quality and the performance of the team that a low-time player participates greatly affects their plus-minus ratings. As a result, poor players on good teams are overrated, while good players on poor teams are underrated.
A common way to resolve the problem is to impose a threshold on the playing time and exclude the corresponding LTPs from the calculation of the performance evaluation metrics; see for example [23] and [17].
Subsequently, in the following, we consider as LTPs all players with playing time less than 200 minutes for the whole of the season. This threshold, selected to be between the recommended thresholds used by [23] and [17], identifies 232 (out of 717) as LTPs which is the $32\%$ of the players in our dataset.
@@ 3.3.2 Linear normal regression with shrinkage methods
for $i=1,\dots,n$; where $n$ is the total number of possessions in our dataset, $K$ is the total number of players under consideration, $Pts_{i}$ is the number of points achieved in possession $i$, $\mu_{i}$ is the expected points for possession $i$ for this model, $X_{ik}^{o}$ and $X_{ik}^{d}$ are the indicators of player $k$ in possession $i$ for the offensive and the defensive team, $X_{ij}$ refers to the extra covariate $j$ for possession $i$, $p$ is the number of extra covariates used in the model formulation, $\beta^{o}_{k}$ and $\beta^{d}_{k}$ are the offensive and defensive, respectively, RAPM coefficients for player $k$, $b_{k}$ are the effects of any extra covariates and $b_{0}$ is the overall constant of the model (usually referring to a reference average player obtained by players with zero or very close to zero RAPM coefficients). Finally, $\sigma^{2}$ is the error (or unexplained) variance which specifies the accuracy of the fitted model. In many published articles [see for example 16, 17], the number of points per possession is multiplied by a factor of 100.
@@
Hence interest lies in the estimation of $\beta_{k}^{o}$ and $\beta_{k}^{d}$ coefficients (i.e. on ORAPM and DRAPM, respectively). The constant term $b_{0}$ here is a nuisance parameter with no direct or meaningful interpretation under the current formulation. The RAPM coefficients $\beta^{o}_{k}$ and $\beta^{d}_{k}$ measure the offensive and defensive, respectively, contribution of player $k$ for $k=1,\ldots,K$ players. Pairwise comparison between two players (denoted by $k$ and $m$) can be achieved by considering the RAPM differences $\beta^{o}_{k}-\beta^{o}_{m}$, $\forall k,m=1,\ldots,K,k\neq m$. These differences represent the mean discrepancy in the expected points scored by each player’s team when player $k$ substitutes player $m$ in a specific possession, with all other players on the court staying the same. Similarly, the difference between two defensive RAPMs $\beta^{d}_{k}-\beta^{d}_{m}$ ($\forall k,m=1,\ldots,K,k\neq m$) expresses the average discrepancy in points of each player’s team conceded (by the opponent team) when players $m$ play instead of player $k$ with all other players on the court staying the same. Hence, the larger the difference is, the higher the offensive and defensive contribution of $k$ player.
@@
The related literature mostly suggests ridge regression to derive RAPM ratings. Ridge regression shrinks coefficients towards zero but does not set them exactly equal to zero. This has the advantage that every player will get a non-zero evaluation coefficient for his offensive and defensive contributions; for more details about the ridge regression see [11] and in [16, 13, 6] for the implementation in obtaining basketball RAPMs evaluation metrics.
In the implementation of ridge regression, for the estimation of RAPM coefficients, we have used 10-fold cross-validation in order to specify the tuning shrinkage parameter of ridge regression (often denoted by $\lambda$). In the ridge bibliography, one of the standard choices is to use $\lambda$ which minimizes the average mean root square error (RMSE) obtained from cross-validation. This value was found to be equal to $\lambda_{min}=241.07$ which is considerably lower than the standard value of 2000 reported in RAPM bibliography [see for example 25]. The shrinkage percentage with the choice of $\lambda_{min}$ was extremely high ($>99\%$).
A clear issue that emerged in the RAMP-based player rankings, is that low-time players (LTPs), namely players with less than 200 minutes played in the entire season, appear at the top of the evaluation list with exceedingly high RAPM ratings. This is in line with the literature; refer, for instance, to in [21]. To address this issue, we opted to implement lasso instead of ridge and, in a second step, to exclude LTPs from the analysis or to eliminate their RAPMs from the regression evaluation metric.
@@
Lasso has the extra property, in comparison to ridge, that shrinks smaller coefficients exactly to zero. This provides a more meaningful interpretation of the intercept coefficient of the implemented model, since now it represents the average contribution of a reference player in each possession. By reference player, we mean any player belonging to the group of players with zero coefficients. A total of 553 offensive and 574 defensive players (out of the 717) have been identified to belong in this reference group. Therefore 184 offensive and 143 defensive players were found with non-zero lasso RAPM ratings. To derive the lasso RAPM ratings we have used the shrinkage parameter value of $\lambda_{min}=0.27$ obtained by a 10-fold cross-validation.